A Complete Bibliography of Publications in Computer Physics Communications: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

(a, b) [SS69]. (a, bγ) [SS69]. (a, bγ − γ) [SS69]. d [All69b]. p [All69b].

addresses [Ano69a]. analysis [BPA69]. angular [SS69]. arbitrary [SS69]. Asymptotic [Nor69]. atoms [BZ69]. authors [Ano69c].

Barbara [SS69]. become [Ano69d]. Board [Ano69b]. bound [Smi69a].

calculate [All69a, FS69]. calculating [BT69]. calculations [BZ69]. closed [BZ69]. coefficients [All69b, FS69]. complex [TR69]. compound [SS69]. Computation [SS69]. Condon [All69a]. configuration [BZ69]. core [BZ69]. Coulomb [TR69]. coupled [Nor69]. cross [SS69].

differential [SS69]. double [SS69]. double-differential [SS69].

Editorial [Ano69a, Ano69b]. electron [Nor69]. electrons [All69b]. energies [TR69]. equations [Nor69]. equivalent [All69b]. factors [All69a]. Feshbach [SS69]. Fluxoids [Kam69]. formalism [SS69]. Fortran [Rob69]. Fractional [All69b]. Franck [All69a]. function [Smi69a]. functions [TR69]. generalized [SS69]. Ginzburg [Kam69]. Hauser [SS69]. having [BZ69]. hydrogenic [FS69]. II [SS69]. Instructions [Ano69c]. interaction [BZ69]. Introduction [Bur69]. ions [FS69]. Landau [Kam69]. library [Ano69d]. Mandy [SS69]. Moldauer [SS69]. momenta [SS69]. Mössbauer [BPA69]. non [BZ69]. non-relativistic [BZ69]. Nuclear [Smi69a, Smi69b, SS69]. orbitals [BZ69]. Parameter [Smi69c]. parentage [All69b]. penetrability [Smi69b]. phase [Smi69b]. potential [BT69]. program [All69a, Ano69d, BPA69, BT69, FS69]. programs [Rob69, SS69]. publication [Rob69]. radiative [FS69]. reactions [SS69]. recombination [FS69]. Regge [BT69]. Relativistic [BZ69]. scattering [BT69, Nor69]. scientific [Rob69]. search [Smi69c]. sections [SS69]. shell [All69b]. shift [Smi69b]. solution [Nor69]. spectra [BPA69]. spin [BZ69]. spin-orbitals [BZ69]. state [Smi69a]. subroutine [Smi69a, Smi69b, Smi69c]. subscriber [Ano69d]. total [SS69]. trajectories [BT69]. two [BZ69]. type [SS69]. valence [BZ69]. Versatile [BPA69]. wave [Smi69a].

References

Allison:1969:PCF

[All69a] A. C. Allison. A program to calculate Franck–Condon factors. Computer Physics Communications, 1(1):21–24, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900149.

Allison:1969:FPC

[All69b] D. C. S. Allison. Fractional parentage coefficients for equivalent p shell and equivalent d shell electrons. Computer Physics Communications, 1(1):15–20, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465569900137.

Anonymous:1969:EA

[Ano69a] Anonymous. Editorial addresses. Computer Physics Communications, 1(1):vii–ix, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465569900095.

Anonymous:1969:EB

[Ano69b] Anonymous. Editorial board. Computer Physics Communications, 1 (1):??, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465569900071.

Anonymous:1969:IA

[Ano69c] Anonymous. Instructions to authors. Computer Physics Communications, 1(1):x–xviii, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465569900101.

Anonymous:1969:PLH

[Ano69d] Anonymous. The program library and how to become a subscriber. Computer Physics Communications, 1(1):64–66, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900198.

Bent:1969:VPA

[BPA69] Michael F. Bent, Börje I. Persson, and David G. Agresti. Versatile program for analysis of Mössbauer spectra. Computer Physics Communications, 1(2):67–87, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465569900010.

Burke:1969:PCR

[BT69] P. G. Burke and C. Tate. A program for calculating Regge trajectories in potential scattering. Computer Physics Communications, 1(2):97– 105, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465569900034.

Burke:1969:I

[Bur69] P. G. Burke. Introduction. Computer Physics Communications, 1(1): v–vi, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465569900083.

Beck:1969:RNR

[BZ69] Donald R. Beck and Richard N. Zare. Relativistic and non-relativistic configuration interaction calculations for atoms having a closed core and two valence spin-orbitals. Computer Physics Communications, 1(2):113–134, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465569900058.

Flower:1969:PCR

[FS69] D. R. Flower and M. J. Seaton. A program to calculate radiative recombination coefficients of hydrogenic ions. Computer Physics Communications, 1(1):31–34, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465569900162.

Kammerer:1969:GLF

[Kam69] U. Kammerer. Ginzburg–Landau fluxoids. Computer Physics Communications, 1(1):10–14, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465569900125.

Norcross:1969:ASC

[Nor69] David W. Norcross. Asymptotic solution of coupled equations for electron scattering. Computer Physics Communications, 1(2):88–96, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465569900022.

Roberts:1969:PSF

[Rob69] K. V. Roberts. The publication of scientific Fortran programs. Computer Physics Communications, 1(1):1–9, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900113.

Smith:1969:NBS

[Smi69a] William R. Smith. Nuclear bound state wave function subroutine. Computer Physics Communications, 1(1):55–63, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900186.

Smith:1969:NPP

[Smi69b] William R. Smith. Nuclear penetrability and phase shift subroutine. Computer Physics Communications, 1(2):106–112, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900046.

Smith:1969:PSS

[Smi69c] William R. Smith. Parameter search subroutine. Computer Physics Communications, 1(2):135–140, October 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046556990006X.

Sheldon:1969:CTD

[SS69] Eric Sheldon and Richard Michael Strang. Computation of total, differential, and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bγ) and (a, bγ − γ) (II) generalized programs “Mandy” and “Barbara” for arbitrary angular momenta in Hauser–Feshbach–Moldauer formalism. Computer Physics Communications, 1(1):35–53, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465569900174.

Tamura:1969:CFC

[TR69] Taro Tamura and Frank Rybick. Coulomb functions for complex energies. Computer Physics Communications, 1(1):25–30, July 1969. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0010465569900150.